a bus moves from rest with a uniform acceleration of 2ms² for the first 10 seconds. it then accelerate at a uniform rate of 1ms² for another 15seconds. it continues at a constant speed for 70seconds and finally came to rest in 20seconds by uniform acceleration. calculate the average retardation as the body is brought to rest


1
Expert's answer
2021-02-03T16:12:50-0500

"v_1=a_1t_1=20~\frac ms."

"s_1=\frac{a_1t_1^2}{2}=100~m."

"v_2=v_1+a_2t_2=20+15=35~\frac ms."

"s_2=v_1t_2+\frac{a_2t_2^2}{2}=300+112.5=412.5~m."

"s_3=v_2t_3=2450~m."

"a_3=\frac{v_2}{t_4}=1.75~\frac{m}{s^2}."

"s_4=\frac{a_3t_4^2}{2}=350~m."

"s=s_1+s_2+s_3+s_4=3312.5~m."

"t=t_1+t_2+t_3+t_4=115~s."

"v=\frac st\approx 28.8~\frac ms."